OpenStudy (anonymous):

how do you find sin^2(135 degrees)? trig final tomorrow! please help :(

zepdrix (zepdrix):

135 degrees is one of your special angles, having trouble remembering that one? :)

OpenStudy (anonymous):

apparently...i really have no clue what to do

OpenStudy (anonymous):

\[\sqrt{2}/2 for \sin\]

OpenStudy (anonymous):

135 degree equal to 3pi/4

OpenStudy (anonymous):

My friend found that it equaled 1/2...and i believe this is correct i just don't understand how he got it. can someone explain?

OpenStudy (anonymous):

it is in this form\[\sin ^{2}(3\pi/4)\]

OpenStudy (anonymous):

the answer with calculator is 0.08

geerky42 (geerky42):

sin²(135º) = sin(135º)·sin(135º)